Using the previous tutorial, address the following: Imagine an experiment where three dice are tossed and the numbers on each die is recorded under Die1, Die2 and Die3. Answer the following questions about the sum of the three numbers recorded from: Die1+Die2+Die3. (Discussions allowed)

Hint:

Simulate this experiment 1000 times. (use the same procedure as in the above tutorial, but for the Number of Variables, instead of 1 put 3; since we are rolling 3 dice not one).

Next, create a new column by writing, in cell D1 =A1+B1+C1, and scroll down the results by clicking on D1 cell and bottom-right corner. (for more detail see the Book chapter on lecture 26)

Finally create the histogram based on this new column. Use bins 3,4,5,…,18.

16. What is the (approximate) probability that Die1+Die2+Die3=8?

a. Probability is approximately 5%

b. Probability is approximately 10%

c. Probability is approximately 20%

d. Probability is approximately 25%

17. What is the (approximate) probability that Die1+Die2+Die3<7?

a. Probability is approximately 5%

b. Probability is approximately 10%

c. Probability is approximately 15%

d. Probability is approximately 20%

18. What is the (approximate) probability that Max(Die1,Die2,Die3)=5? Careful: Use “Max( ) function in Excel.

a. Probability is approximately 5%

b. Probability is approximately 15%

c. Probability is approximately 25%

d. Probability is approximately 45%
Answer #1

16)from simulation 8 as a sum appear 113 times therefore approximate probabilty =113/1000~ 10%

option B is correct

17)

sum is <7; 95 times;l hence probability =95/1000 ~10% ; option B

18)

5 as maximum appears 290 times ; hence probability=290/1000 ~0.29

option C